Session: Jun 2012 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Thu May 31 11:05:17 2012) Minor failures: RECO
We lose 7 scans from day 150 08:50UT to day 150 09:43UT due to recorder problem at bank switch

Hartebeesthoek (Jonathan Quick at Thu May 31 12:31:38 2012) Minor failures:RECE
Maser failed during experiment at 149/08:00, new maser installed at ~149/08:40, data will be invalid for the 5 scans in that period.

Yebes (Pablo de Vicente at Sat Jun 02 07:04:05 2012) Minor failures: RECO
This experiment had the wrong patching. We changed it at 14:00 day May28th. IFA-> 8150 MHz, B-> 7650 MHz C -> S band Subscans 072, 0713 and 0717,were lost Mak5B had problems from 07:26:06 UT to 10.57:55 UT. Lost scans from 1348 to 1484(include).

Onsala (Michael Lindqvist at Sun Jun 03 04:16:55 2012) Minor failures:
We did not have any phasecal first night. Replaced in the morning. In addition, we had problems with IF3. We also had problems with the Mark5 (first night). Some bad weather, rain.

Noto (P. Cassaro at Tue Jun 05 15:26:23 2012) Minor failures: TSYS
Tsys are missing during the whole experiment.

Zelenchukskaya (Kononenko V. and Zabavskaya N. at Thu Jun 07 12:13:30 2012) Success:

Effelsberg (Uwe Bach at Wed Jun 13 08:58:20 2012) Minor failures: PCAL
2012.150.08:14:20.33;"phascal was off until now 2012.151.12:26:17.92;"off source since 12:20 due to a technical problem 2012.151.12:50:48.39;"back on source, 12:50. Emergency stop was hit by accident A short stop because of a thunderstorm occurred on DOY 150, 21:10 UT. Unfortunately 50% of the experiment was done with phase-cal off, because of a software failure.